OpenStudy (anonymous):

Simplify the expression 3x(2x^2).

sam (.sam.):

6x^3

OpenStudy (callisto):

3x(2x^2) = 3(2) x(x^2) = 6 ( x^(1+2) ) =?

OpenStudy (anonymous):

6x^3???????

OpenStudy (callisto):

Yes